A pair of dice is rolled, and a number that appears uppermost on each die is observed, Refer to this experiment and find the probability of the given event. (Enter your answer as a fraction)The sume of the numbers ia an add number.
Answer #1
Would it not make logical sense that half the sums would be even and half would be odd ?

e.g.
1,1 -- even
1,2 --odd
1,3 -- even
...
6,1 -- odd
6,2 -- even
..
6,6, -- even , count them if that doesn't make sense

prob(odd sum) = 18/36 = 1/2
answered by: cally